OpenStudy (anonymous):

determine whether this series converges or diverges sigma 1 to infinity -1^(n-1)2^n/(n^2)

OpenStudy (anonymous):

These fractional types usually suggest ratio test

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{(-1)^{n-1} 2^n}{n^2}\] as chag said, using the ratio test, you'll get: \[|\frac{an+1}{an}| = |\frac{(-1)^n2^{n+1}}{(n+1)^2} . \frac{n^2}{(-1)^{n-1}2^n}|\]

OpenStudy (anonymous):

so :

OpenStudy (anonymous):

\[= \frac{2^n.2^1}{(n+1)^2} . \frac{n^2}{2^n} = \frac{2n^2}{(n+1)^2}\] now the second step is to find the limit as n--> infinity: \[\lim_{n \rightarrow \infty} \frac{2n^2}{(n+1)^2} = 2\] and the Ratio Test's theorem says the following: if L (limit) > 1 = series diverge if L <1 = series converge if L = 1 , no conclusion Correct me if I'm wrong please ^_^

OpenStudy (anonymous):

so for this case, since 2 > 1, then the following series diverge :)

OpenStudy (anonymous):

if you don't understand what I did, let me know :)

OpenStudy (anonymous):

ty i got it

OpenStudy (anonymous):

np ^_^

OpenStudy (anonymous):

can u help me with another problem

OpenStudy (anonymous):

I'll give it a try ^_^